In deriving the LM curve, we assumed that money demand was negatively related to the interest rate. Suppose instead that money demand depends only on the real value of transactions, Y . Draw the money demand curve for this case. What does the LM curve look like if money demand has this property? Explain. (Note: If money demand is independent of the interest rate, then we have the quantity theory of money.
Answer #1

Hi! Welcome to Chegg!

Ans:

General form of money demand:

Md = kY - hi

where,

Y = real output

i = interest rate

k, h are constant parameters.

When money demand is independent of interest rate (i), we have:

Md = kY

Money market equilibrium (LM curve):

Money supply = Money demand

M = kPY (LM curve)

This corresponds to quantity theory of money:

M/k = PY

Since, both money supply and demand are independent of interest rate, LM curve is also vertical.

If you are satisfied with the answer, please provide a positive rating. Feel free to comment in case of queries.

Have a nice day ahead!